# Differentially Private Matchings

### v2

@@ 5 $O(\log n)$-Round $\varepsilon$-LEDP Implicit Matchings
See 1.6
High-Level Overview
@@ 5.1 Detailed Algorithm Description
We now describe our pseudocode for Algorithm 3 in detail and then prove its properties. We use a number of constants that guarantees our $\Theta(\log n)$ round algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$. Thus, in our pseudocode, some of our constants depend on $c$. Our algorithm returns a maximal matching with high probability when we set $b=\Omega(\log^{2}(n)/\varepsilon)$.
In our pseudocode given in Algorithm 3, we define $M(u)$ to be the set of nodes currently matched to $u\in V$. For each round $i$, we let $V_{i}$ be the set of active vertices which have not reached their matching capacity in round $i$. Our algorithm uses a subroutine (Algorithm 4) which takes a vertex $v$, a set of vertices $S$ from which to select a proposal or matching set, and returns a subgraph index which determines a subset of vertices based on public i.i.d. coin flips. This algorithm implements a more specific version of our Public Vertex Subset Mechanism.
Algorithm 3 takes as input a graph $G=(V,E)$, a privacy parameter $\varepsilon>0$, and a matching parameter $b$ where $b=\Omega(\log^{2}(n)/\varepsilon)$. The algorithm returns an $\varepsilon$-LEDP implicit $b$-matching. First, we set some additional approximation and privacy parameters in Algorithm 3. We then iterate through every node in Algorithm 3 simultaneously to determine the noisy threshold of every node (Algorithm 3). This threshold is used to determine if a vertex satisfies its matching condition. In particular, if the estimated number of possible matches for the node is greater than the threshold, then with high probability, the node is matched to at least one neighbor. The threshold is set in Algorithm 3 using Laplace noise. We add Laplace noise to the threshold as part of an instance of the Multidimensional AboveThreshold (MAT) technique (Lemma 2.15). This is in turn used to determine for every node whether the number of matches exceeds this threshold. We also initialize an empty set for each $u\in V$, denoted $M(u)$, that contains the set of nodes $v$ is matched to (Algorithm 3).
@@
We then flip another set of coins to determine which nodes are proposers (Algorithm 3) in round $i$. Each node is a proposer with $1/2$ probability; the result of the coin flip is stored in $a_{i}(v)$ for each $v\in V$. We only select proposers from the set of active vertices, $V_{i}$. For each selected proposer, we simultaneously call the procedure PrivateSubgraph with the inputs $G,\varepsilon^{\prime},\eta,w,V_{i}\setminus P_{i}$ and $coin$ (Algorithms 3 and 3). The pseudocode for PrivateSubgraph is given in Algorithm 4. The function takes as input a graph $G$, a privacy parameter $\varepsilon^{\prime}$, an approximation parameter $\eta$, a vertex subset $S$, a vertex $w\in V$, and all public coin flips $coin$. The function then iterates through all possible subgraph indices $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$ (Algorithm 4). For each index, we determine the set of nodes $u$ that satisfy the following conditions: $u\in S$, the edge $\{w,u\}$ exists, and the coin flip $coin(u,w,r)$ is Heads. Since the set $W_{r}(w)$ is computed privately and not released, $w$ has access to the information about whether $\{w,u\}\in E$. This set of nodes is labeled $W_{r}(w)$ (Algorithm 4). We then add Laplace noise to the size of this set to obtain a noisy estimate for the size of $W_{r}(w)$ (Algorithm 4). We use the Adaptive Laplace Mechanism to determine the smallest $r$ (most number of nodes $w$ can propose to) that does not exceed $147c\log(n)/\varepsilon^{\prime}$. To do this, we add Laplace noise to the size of the set of nodes matched to $w$ (Algorithm 4) and find the smallest $r$ such that the sum of the noisy proposal set size, $|\widetilde{W}_{r}(w)|$, and the noisy matched set size, $|\widetilde{M}(w)|$, plus $27c\log(n)/\varepsilon^{\prime}$ does not exceed $147c\log(n)/\varepsilon^{\prime}$ (Algorithm 4). The term $27c\log(n)/\varepsilon^{\prime}$ is added to ensure that $|W_{r}(w)|+|M(w)|$ does not exceed our $b$ bound (by drawing negative noises), with high probability. The procedure releases the smallest subgraph index (Algorithm 4) satisfying Algorithm 4.
@@
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, matching parameter $b\geq 147c\log^{2}(n)/\varepsilon$
@@
2Let $\eta\leftarrow 1/2,\varepsilon^{\prime}\leftarrow\varepsilon/(192c\log_{16/15}(n)),\varepsilon^{\prime\prime}\leftarrow\varepsilon/3$
@@
       $\tilde{b}(u)\leftarrow 38c\log_{1+\eta}(n)/\varepsilon^{\prime\prime}+\text{Lap}(4/\varepsilon^{\prime\prime})$ // Noisy threshold
@@ 4       $M(u)\leftarrow\varnothing$
7 for round $i=1$ to $\lceil 32c\cdot\log_{16/15}(n)\rceil$ do
       // MAT for checking matching capacity
@@
9             $\nu_{i}(u)\leftarrow\text{Lap}(8/\varepsilon^{\prime\prime})$
@@ 11                   Release: node $u$ has satisfied matching condition and remove $u$ from $V_{i}$
16
      // Curator flips coins for proposal and match sets; proposal set coins are stored in $coin_{p}$ and matching set coins in $coin_{m}$
@@
             // If $v$ is a proposer, find the proposal set
             $r_{w}\leftarrow$ PrivateSubgraph$(G,\varepsilon^{\prime},\eta,w,V_{i}\setminus P_{i},coin_{p})$   // Algorithm 4
@@ 29             $R\leftarrow\{u\mid w\in W_{r_{u}}(u)\wedge u\in P_{i}\}$
             // Determine the match set
             $r_{w}\leftarrow$ PrivateSubgraph$(G,\varepsilon^{\prime},\eta,w,R,coin_{m})$   // Algorithm 4
@@ 40      $V_{i+1}\leftarrow V_{i}$
Algorithm 3 $O(\log n)$-Round $\varepsilon$-LEDP Maximal Matching
Input: Graph $G=(V,E)$, privacy parameters $\varepsilon^{\prime}$, approximation parameter $\eta$, vertex set $S$, vertex $v\in V$, and public coin flips $coin$
@@
2Function PrivateSubgraph($G,\varepsilon^{\prime},\eta,w,S,coin$) begin
@@
             // Determine sets with exponentially increasing size
@@ 4             $W_{r}(w)\leftarrow\{u\mid u\in S\wedge\{w,u\}\in E\wedge coin(u,w,r)=\textsc{Heads}\}$
      $|\widetilde{M}_{i}(w)|\leftarrow|M(w)|+\text{Lap}(2/\varepsilon^{\prime})$ // Noisy count of matched nodes
       $r\leftarrow\min\left(\{r:|\widetilde{M}_{i}(w)|+|\widetilde{W}_{r}(w)|+27c\log(n)/\varepsilon^{\prime}\leq 147c\log(n)/\varepsilon^{\prime}\}\right)$ // Smallest $r$
@@ 5.2.1 Privacy Guarantees
Our privacy proof follows a similar flavor to the $b$-matching privacy proof given in the previous section except for one main difference. While the adaptive Laplace mechanism was called twice per node in its use in the previous algorithm, this is not the case in our distributed algorithm. Suppose that given neighbor graphs $G$ and $G^{\prime}$ with edge $\{u,v\}$ that differs between the two. Nodes $u$ and/or $v$ can propose a set many times during the course of the algorithm since their proposed sets are not guaranteed to be matched. If they are particularly unlucky, they could propose a set during every round of the algorithm. This means that by composition, we privacy loss proportional to the number of rounds. We formally prove the privacy of our algorithm below.
@@
To prove that Algorithm 3 is $\varepsilon$-LEDP, we show that our algorithm is implemented using local randomizers. Each node $v$ in the algorithm only releases the following set of information in Algorithms 3, 3 and 3: when $v$ has satisfied its matching condition (Algorithm 3), when $v$ releases a proposal subgraph index (Algorithm 3), and when $v$ releases a match set index (Algorithm 3).
@@
Each node implements three different local randomizers for releasing each of the aforementioned three types of information. Our local randomizers are created for each type of private information used to determine the released output. Namely, we use a local randomizer to determine when a node has satisfied the matching condition (Algorithm 3), a local randomizer for computing the noisy proposal or match sets in Algorithm 4 of Algorithm 4, and a local randomizer for determining the noisy set of matched edges in Algorithm 4. The releases in Algorithms 3 and 3 solely depend on the noisy proposal/match sets and the noisy set of matched edges. Then by the concurrent composition theorem, the entire algorithm is differentially private since no other parts of the algorithm use the private graph information.
@@
First, before we dive into the privacy components, we note that the curator flip coins and releases the result of the coins (Algorithms 3 and 3); these coin flips do not lose any privacy since the coins are not tied to private information. The coin flips are performed for each public pair of distinct nodes. In other words, these coins only use the set of nodes which is public. Our proof follows the privacy proof of Lemma 4.1 except we need to account for multiple uses of the adaptive Laplace mechanism for each node.
@@
The first set of local randomizers for all nodes $u\in V$ that releases whether $u$ has satisfied its matching condition can be implemented using an instance of the Multidimensional AboveThreshold (MAT) mechanism with sensitivity. We show how Algorithms 3, 3, 3, 3 and 3 can be implemented using MAT. First, our vectors of queries at each round $i$ is an $n$-length vector which contains the number of nodes each node $u\in V$ has already matched to. Specifically, the vector of queries at each round $i$ is the number of other nodes, $|M(u)|$, each node $u$ has already been matched to before round $i$. Conditioning on the outputs of the previous iterations, the addition or removal of a single edge can only affect $M(u)$ for two nodes, each by at most $1$. Hence, the sensitivity of the vector of queries is $2$; furthermore, the sensitivity of each vector of queries for all rounds is $2$. This means that MAT can be implemented with $\Delta_{M}=2$. Thus, the first set of local randomizers is $\varepsilon^{\prime\prime}$-differentially private by Lemma 2.15.
@@
The second type of local randomizer computes the $|\widetilde{W}_{r}(w)|$ values for each $w\in V$ that is a proposer or receiver. As before, each call of Algorithm 4 can be implemented as a local randomizer using the Adaptive Laplace Mechanism with sensitivity $\Delta=1$; the sensitivity is $1$ because on neighboring adjacency lists, $|W_{r}(w)|$ differs by at most $1$. In fact, given that each edge is selected with probability $p_{r}=(1+\eta)^{-r}$, we can give the privacy guarantee for this local randomizer in terms of $p$ using Lemma 2.17. For a given randomizer and parameter $r$, by Lemma 2.17, the randomizer is $\frac{2}{(1+\eta)^{r}}\cdot\frac{\varepsilon^{\prime}}{4}=\frac{\varepsilon^{\prime}}{2(1+\eta)^{r}}$-DP. Then, the sum of the privacy parameters of all calls to this second type of local randomizer on node $w$ is
$\sum_{r=0}^{\lceil\log_{1+\eta}(n)\rceil}\frac{\varepsilon^{\prime}}{2(1+\eta)^{r}}=\frac{\varepsilon(\eta n+n-1)}{2\eta\cdot n}\leq\frac{\varepsilon^{\prime}\cdot(\eta+1)}{2\eta}=\frac{3\varepsilon^{\prime}}{2}.$
@@
$2\cdot\frac{3\varepsilon^{\prime}}{2}\cdot 32c\log_{16/15}(n)=3\cdot\frac{\varepsilon}{192c\log_{16/15}(n)}\cdot 32c\log_{16/15}(n)=\frac{\varepsilon}{2}\,.$
@@
$2\cdot\frac{\varepsilon^{\prime}}{2}\cdot 32c\log_{1+\eta}(n)=2\cdot\frac{\varepsilon}{2\cdot 192c\log_{1+\eta}(n)}\cdot 32c\log_{1+\eta}(n)=\frac{\varepsilon}{6}\,.$

### v3

@@ 6 $O(\log n)$-Round LEDP Implicit $b$-Matching via PVSM
See 2.3
As an immediately corollary for the case $b^{\prime}=1$, we obtain the following.
Corollary 6.1.
Let $\varepsilon\in(0,1)$ and $b=\Omega\left(\frac{\log^{2}(n)}{\varepsilon}\right)$. Algorithm 3 is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an (implicit) $b$-matching in the billboard model whose size is at least a $2$-approximate maximum $1$-matching.
Similar to the sequential algorithm, for $b$ sufficiently large, we can ensure an approximation guarantee with respect to the maximum $b$-matching.
Theorem 6.2.
Let $\varepsilon\in(0,1)$ and $b=\Omega\left(\frac{\log^{2}(n)}{{\eta^{5}}\varepsilon}\right)$. Algorithm 3 is an $\varepsilon$-LEDP algorithm that terminates in $O(\log n)$ rounds and, with high probability, outputs an (implicit) $b$-matching in the billboard model whose size is at least a $(2+\eta)$-approximate maximum $b$-matching.
High-Level Overview.
@@ 6.1 Detailed Algorithm Description
We now describe our pseudocode for Algorithm 3 in detail and then prove its properties. We use a number of constants that guarantees our $\Theta(\log n)$ round algorithm succeeds with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$. Our algorithm returns a $(2+\eta)$-approximate $b^{\prime}$-matching with high probability when the degree parameter $b$ satisfies the bound given in Theorem 2.3.
In our pseudocode given in Algorithm 3, we define $M(u)$ to be the set of nodes currently matched to $u\in V$. For each round $i$, we let $V_{i}$ be the set of active vertices which have not reached their matching $b^{\prime}$-condition in round $i$. Our algorithm uses a subroutine (Algorithm 4) which takes a vertex $v$, a set of vertices $S$ from which to select a proposal or matching set, and returns a subgraph index which determines a subset of vertices based on public i.i.d. coin flips. This algorithm implements a more specific version of our Public Vertex Subset Mechanism.
Algorithm 3 takes as input a graph $G=(V,E)$, a privacy parameter $\varepsilon>0$, a matching benchmark parameter $b^{\prime}\geq 1$, and a degree parameter $b$. The algorithm returns an $\varepsilon$-LEDP implicit $b$-matching. First, we set some additional approximation and privacy parameters in Algorithm 3. We then iterate through every node in Algorithm 3 simultaneously to determine the noisy threshold of every node (Algorithm 3). This threshold is used to determine if a vertex satisfies its matching condition. In particular, if the estimated number of possible matches for the node is greater than the threshold, then with high probability, the node is matched to at least $b^{\prime}$ neighbors. The threshold is set in Algorithm 3 using Laplace noise. We add Laplace noise to the threshold as part of an instance of the Multidimensional AboveThreshold (MAT) technique (Lemma 3.16). This is in turn used to determine for every node whether the number of matches exceeds this threshold. We also initialize an empty set for each $u\in V$, denoted $M(u)$, that contains the set of nodes $v$ is matched to (Algorithm 3).
@@
We then flip another set of coins to determine which nodes are proposers (Algorithm 3) in round $i$. Each node is a proposer with $1/2$ probability; the result of the coin flip is stored in $a_{i}(v)$ for each $v\in V$. We only select proposers from the set of active vertices, $V_{i}$. For each selected proposer, we simultaneously call the procedure PrivateSubgraph with the inputs $G,\varepsilon^{\prime},\eta,b,$ $w,V_{i}\setminus P_{i}$ and coincoin (Algorithms 3 and 3). The pseudocode for PrivateSubgraph is given in Algorithm 4. The function takes as input a graph $G$, a privacy parameter $\varepsilon^{\prime}$, an approximation parameter $\eta$, the degree cap $b$, a vertex subset $S$, a vertex $w\in V$, and all public coin flips coincoin. The function then iterates through all possible subgraph indices $r\in\{0,\dots,\lceil\log_{1+\eta}(n)\rceil\}$ (Algorithm 4). For each index, we determine the set of nodes $u$ that satisfy the following conditions: $u\in S$, the edge $\{w,u\}$ exists, and the coin flip $coin(u,w,r)$ is Heads. This set of nodes is labeled $W_{r}(w)$ (Algorithm 4). We then add Laplace noise to the size of this set to obtain a noisy estimate for the size of $W_{r}(w)$ (Algorithm 4). We use the Adaptive Laplace Mechanism to determine the smallest $r$ (most number of nodes $w$ can propose to) that does not exceed $b$. To do this, we add Laplace noise to the size of the set of nodes matched to $w$ (Algorithm 4) and find the smallest $r$ such that the sum of the noisy proposal set size, $|\widetilde{W}_{r}(w)|$, and the noisy matched set size, $|\widetilde{M}(w)|$, plus $27c\log(n)/\varepsilon^{\prime}$ does not exceed $b$ (Algorithm 4). The term $27c\log(n)/\varepsilon^{\prime}$ is added to ensure that $|W_{r}(w)|+|M(w)|$ does not exceed our $b$ bound (by drawing negative noises), with high probability. The procedure releases the smallest subgraph index (Algorithm 4) satisfying Algorithm 4.
@@
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$, matching benchmark parameter $b^{\prime}\geq 1$, constant $c\geq 1$
@@
2Let $\eta\leftarrow 1/2,\varepsilon^{\prime}\leftarrow\varepsilon/(3072c\log_{16/15}(n)),\varepsilon^{\prime\prime}\leftarrow\varepsilon/3$
3 Set parameter $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+\frac{518\log_{1+\eta}(n)}{\eta^{4}\,\varepsilon^{\prime}}$
@@
5 $\tilde{b}(u)\leftarrow b-259c\log_{1+\eta}(n)/\varepsilon^{\prime}+\mathrm{Lap}(4/\varepsilon^{\prime\prime})$
@@ 6 $M(u)\leftarrow\varnothing$
9 for round $i=1$ to $\lceil 512c\cdot\log_{16/15}(n)\rceil$ do
@@
11    $\nu_{i}(u)\leftarrow\mathrm{Lap}(8/\varepsilon^{\prime\prime})$
@@ 13       Release: node $u$ has satisfied matching condition and remove $u$ from $V_{i}$
18
@@
[…]
@@ 32    $R\leftarrow\{u\mid w\in W_{r_{u}}(u)\wedge u\in P_{i}\}$
[…]
@@ 44 $V_{i+1}\leftarrow V_{i}$
Algorithm 3 $O(\log n)$-Round $\varepsilon$-LEDP Approximate Maximum $b^{\prime}$-Matching
Input: Graph $G=(V,E)$, privacy parameters $\varepsilon^{\prime}$, approximation parameter $\eta$, degree cap $b$, vertex set $S$, vertex $v\in V$, and public coin flips coincoin
@@
2FunctionPrivateSubgraph($G,\varepsilon^{\prime},\eta,b,w,S,coin$) begin
@@ 4    $W_{r}(w)\leftarrow\{u\mid u\in S\wedge\{w,u\}\in E\wedge coin(u,w,r)=\textsc{Heads}\}$
7 $|\widetilde{M}_{i}(w)|\leftarrow|M(w)|+\text{Lap}(2/\varepsilon^{\prime})$
8 $r\leftarrow\min\left(\{r:|\widetilde{M}_{i}(w)|+|\widetilde{W}_{r}(w)|+27c\log(n)/\varepsilon^{\prime}\leq b\}\right)$
@@ 6.2.1 Privacy Guarantees
Our privacy proof follows a similar flavor to the $b$-matching privacy proof given in the previous section except for one main difference. While the adaptive Laplace mechanism was called twice per node in its use in the previous algorithm, this is not the case in our distributed algorithm. Suppose we are given neighbor graphs $G$ and $G^{\prime}$ with edge $\{u,v\}$ that differs between the two. Nodes $u$ and/or $v$ can propose a set many times during the course of the algorithm since their proposed sets are not guaranteed to be matched. If they are particularly unlucky, they could propose a set during every round of the algorithm. This means that by composition, we have privacy loss proportional to the number of rounds. We formally prove the privacy of our algorithm below.
@@
To prove that Algorithm 3 is $\varepsilon$-LEDP, we show that our algorithm can be implemented using local randomizers. Each node $v$ in the algorithm only releases a privatized set of information in the following Algorithms 3, 3 and 3: when $v$ has satisfied its $b^{\prime}$-matching condition (Algorithm 3), when $v$ releases a proposal subgraph index (Algorithm 3), and when $v$ releases a match set index (Algorithm 3).
@@
Each node implements three different local randomizers for releasing each of the aforementioned three types of information. Our local randomizers are created for each type of private information used to determine the released output. Namely, we use a local randomizer to determine when a node has satisfied the $b^{\prime}$-matching condition (Algorithm 3), a local randomizer for computing the noisy proposal or match sets in Algorithm 4 of Algorithm 4, and a local randomizer for determining the noisy set of matched edges in Algorithm 4 of Algorithm 4. The releases in Algorithms 3 and 3 solely depend on the noisy proposal/match sets and the noisy set of matched edges. Then, by the concurrent composition theorem, the entire algorithm is differentially private since no other parts of the algorithm use the private graph information.
@@
First, before we dive into the privacy components, we note that the curator flips coins and releases the result of the coins (Algorithms 3 and 3); these coin flips do not lose any privacy since the coins are not tied to private information. The coin flips are performed for each public pair of distinct nodes. Our proof follows the privacy proof of Lemma 5.4 except we need to account for multiple uses of the adaptive Laplace mechanism for each node.
@@
The first set of local randomizers for all nodes $u\in V$ that releases whether $u$ has satisfied its matching condition can be implemented using an instance of the Multidimensional AboveThreshold (MAT) mechanism with sensitivity $2$. We show how Algorithms 3, 3, 3, 3, 3 and 3 can be implemented using MAT. First, our vectors of queries at each round $i$ is an $n$-length vector which contains the number of nodes each node $u\in V$ has already matched to. Specifically, the vector of queries at each round $i$ is the number of other nodes, $|M(u)|$, each node $u$ has already been matched to before round $i$. Conditioning on the outputs of the previous iterations, the addition or removal of a single edge can only affect $M(u)$ for two nodes, each by at most $1$. Hence, the sensitivity of the vector of queries is $2$; furthermore, the sensitivity of each vector of queries for each of the rounds is $2$. This means that MAT can be implemented with $\Delta_{M}=2$. Thus, the first set of local randomizers is $\varepsilon^{\prime\prime}$-differentially private by Lemma 3.16.
@@
The second type of local randomizer computes the $|\widetilde{W}_{r}(w)|$ values for each $w\in V$ that is a proposer or receiver. As before, each call of Algorithm 4 can be implemented as a local randomizer using the Adaptive Laplace Mechanism with sensitivity $\Delta=1$; the sensitivity is $1$ because on neighboring adjacency lists, $|W_{r}(w)|$ differs by at most $1$. In fact, given that each edge is selected with probability $p_{r}=(1+\eta)^{-r}$, we can amplify the privacy guarantee for this local randomizer in terms of $p$ using Lemma 3.18. For a given randomizer and parameter $r$, by Lemma 3.18, the randomizer is $\frac{2}{(1+\eta)^{r}}\cdot\frac{\varepsilon^{\prime}}{4}=\frac{\varepsilon^{\prime}}{2(1+\eta)^{r}}$-DP. Then, the sum of the privacy parameters of all calls to this second type of local randomizer on node $w$ is
$\sum_{r=0}^{\lceil\log_{1+\eta}(n)\rceil}\frac{\varepsilon^{\prime}}{2(1+\eta)^{r}}=\frac{\varepsilon^{\prime}(\eta n+n-1)}{2\eta\cdot n}\leq\frac{\varepsilon^{\prime}\cdot(\eta+1)}{2\eta}=\frac{3\varepsilon^{\prime}}{2}.$
@@
$2\cdot\frac{3\varepsilon^{\prime}}{2}\cdot 512c\log_{16/15}(n)=3\cdot\frac{\varepsilon}{3072\log_{16/15}(n)}\cdot 512c\log_{16/15}(n)=\frac{\varepsilon}{2}\,.$
@@
$2\cdot\frac{\varepsilon^{\prime}}{2}\cdot 512c\log_{1+\eta}(n)=2\cdot\frac{\varepsilon}{2\cdot 3072c\log_{16/15}(n)}\cdot 512c\log_{1+\eta}(n)=\frac{\varepsilon}{6}\,.$